The half-life of the Winease catalyzed conversion is 10 sec. if you start with 100mM of lemonade, approximately how much lemonade will remain after 100 sec? A.- 1uM B.-10uM C.-100uM D.-10 mM
Answer #1

Half life = 10 sec

rate constant,k =0.693/t = 0..693/10 = 0.0693 sec-1

we know,

0.0693 * 100 = ln (100*10-3)/[A]

or, [A] = 9.78 *10-5

[A] = 100M